OpenStudy (anonymous):

Austin1617 help me with this please. 4 2/3 divided by 1 1/4.

OpenStudy (anonymous):

@Austin1617

OpenStudy (anonymous):

3.73

OpenStudy (chealyn98):

3.73

OpenStudy (anonymous):

how do u work it out

OpenStudy (anonymous):

Exact Form: 56/15 Decimal Form:3.73333 Mixed Number Form: 3 11/15

OpenStudy (knightofpentacles):

i ill tell you the process for dividing fractions

OpenStudy (anonymous):

ok

OpenStudy (knightofpentacles):

1 2/4 2 4/6 is the example to tackle dividing fraction

OpenStudy (anonymous):

\[\frac{ 14 }{ 3\div1, \frac{ 1 }{ 4}} \[\frac{ 14 }{ 3\div1, (\frac{ 5 }{ 4 }) }\] \[\frac{ 14 }{ 3\div (\frac{ 5 }{ 4 }) }\] \[(\frac{ 14 }{ 3 })(\frac{ 4 }{ 5 })\]

OpenStudy (anonymous):

@Cammie19

OpenStudy (knightofpentacles):